# combined_file.bib Non-classical Temperature Dependence of Chirality-Induced Magnetization and Its Implications for RNA’s Homochirality

### v1

@@ Conclusion
[…]
Acknowledgments
We thank John Sutherland for his help in interpreting the NMR spectra and Ziwei Liu for his suggestions on quantitative NMR analysis. We thank Naama Goren from the Hebrew University Applied Physics Department for her help in interpreting the Hall measurements and Maurice Saidian from the Hebrew University Nano Center for Ni/Au evaporation. Y.P. acknowledges the funding from Marie Skłodowska-Curie Actions under Horizon Europe framework program (CISSE project No. 101071886), Carl Zeiss Stiftung (HYMMS project No. P2022-03-044), and U.S Air Force (grant No. FA8655-24-1-7390). Y.K. thanks the Israel Council of Higher Education’s VTT fellowship for women in STEM. S.F.O. and D.D.S. acknowledge the Harvard Origins of Life Initiative for funding and its members for fruitful discussions. S.F.O. also acknowledges the Kavli-Laukien fellowship program and the Kavli and Laukien Foundations for generous research funding and travel support.
@@ Methods
Hall measurements
Hall measurements in Van der Pauw configuration were performed on Ni(40 nm)/Au(10 nm) samples coated with D-RAO molecules. The measurements were conducted under a constant external magnetic field of -25 mT generated by an out-of-plane electromagnet powered by a Kepco BOP 100-4DL power supply. The current was supplied by Keysight B2901A precision source while Hall voltage was measured in parallel by Keitley 2400 SourceMeter. Temperature was controlled by a Peltier element and was measured continuously using a Pt thermistor. The current was scanned from -1 mA to 1 mA and back and forth. These measurements were repeated 5 times for each direction and at different temperatures: 30°C, 40°C, 50°C, 60°C, 70°C, and 80°C. The Hall resistance () was extracted by plotting the Hall voltage as a function of the current and determining the slope from a linear fit to the data, then averaging over different repetitions.
Quantitative NMR measurements
A 1:1 mixture of DL-glyceraldehyde (90 g) and 2-AO (84 g), each at 250 mM, was prepared in 4 mL of DI water. The pH was not adjusted after mixing the reagents. This solution was divided into five tubes, each containing 800 µL of liquid, and incubated at five temperatures: 5°C, 20°C, 35°C, 50°C, and 65°C. Samples were analyzed at two time points: 2.5 hours and 24 hours.
For each NMR measurement, 300 µL of liquid from the sample was extracted and placed in an NMR tube. The NMR tubes were immediately placed in a water ice bath to slow the reaction, ensuring that measurement time would not impact the comparison across samples. 1H NMR spectra were measured in water with a Bruker 400 MHz NMR spectrometer, using the water suppression feature.
[…]
@@ 1.1 MOKE Setup
1.2 Effect of RAO Chirality
The MOKE experiment was repeated using both L- and D-chiralities of RAO crystals. In Figure S6a, the magnetic hysteresis loops of Ni are presented for regions far from (top) and directly beneath (bottom) the D-RAO crystals. For the L-RAO sample, the magnetic properties of the bare Ni surface remained unchanged at all temperatures. In contrast, for the D-RAO sample, the coercive field of the bare Ni surface decreased at higher temperatures. Nonetheless, the difference in the coercive field between the regions beneath and far from the crystals increased with temperature, as shown in Figure S6b. In Figure S6c, this difference is compared for both L- and D-chiralities, where a similar linear trend is observed for both.
@@ 1.7 MOKE Analysis Code
2 Methods for Hall Measurements
2.1 Hall Setup
@@ 2.2 Analysis of Hall Measurements
The Hall Van der Pauw measurements were performed on both bare Ni and Ni samples after drop-casting RAO molecules. The measurements were conducted under varying external magnetic fields of -50 mT, -25 mT, 0 mT, 25 mT, and 50 mT. The out-of-plane magnetic field was generated by an electromagnet obtained from the MOKE microscope supplier powered by a Kepco BOP 100-4DL power supply. For each constant external field, the current was scanned from -1 mA to 1 mA in 101 steps, and the Hall voltage was measured. Each scan was repeated 5 times in the forward direction (-1 mA to 1 mA) and 5 in the reverse direction (1 mA to -1 mA). These measurements were repeated at six different temperatures: 30°C, 40°C, 50°C, 60°C, 70°C, and 80°C. A linear relationship was observed for each measurement of the Hall voltage as a function of the current. A linear function was fitted to each measurement with the slope of the fit being the hall resistance,. For example, the five repetitions in the forward direction of the Hall IV curve and the corresponding fits for Ni/Au with RAO crystals at 50mT and 80°C shown in Figure S12.
was averaged across all repetitions for each specific external field, temperature, and direction (forward or reverse), both with and without the D-RAO molecules. For each of these averages, a standard error was calculated. for all the temperatures and fields are represented in Figure S13.
As shown, the behavior is similar for all magnetic fields, even for the negative ones. We expected to observe a negative Hall voltage for the negative fields, but this was not the case. This, along with the relatively small external fields applied (up to 50 mT), suggests that the effect we measured is the Anomalous Hall Effect, and in the applied field range, the magnetization of the Ni did not change. The -25 mT field was chosen for Figure 3 arbitrarily. with the molecules is much larger compared to the bare Ni without molecules and compared to the temperature effect. We, therefore, subtracted taken at 30°C from all data points. This way, both the D-RAO and bare substrate results start at zero, and the temperature increase can be observed as seen in Figure S13b.
@@ 2.3 Hall Analysis Code
3 Methods for NMR Measurements
The prebiotic synthesis of pentose-aminooxazolines from glyceraldehyde and 2-aminooxazole (2-AO) yields a mixture of four stereoisomers and their corresponding enantiomers: ribose-, arabinose-, xylose-, and lyxose-aminooxazoline.
Among these, the ribose form (RAO) is the target stereoisomer, which can be obtained in enantiopure form and serves as a direct precursor to RNA pyrimidines and DNA purines. The relative abundance of RAO in the mixture increases with temperature, as determined by 1H NMR spectroscopy in water.
Samples containing a mixture of pentose aminooxazolines in water were incubated at five temperatures (5°C, 20°C, 35°C, 50°C, and 65°C) and probed at two time intervals: 2.5 hours and 24 hours.
3.1 Quantitative NMR Analysis
NMR analyses were conducted on the H (C1’) doublets of pentose aminooxazolines using a Bruker 400 MHz NMR spectrometer, with Mnova software employed for spectrum processing and integration.
[…]
The 0 ppm TSP-d4 singlet was employed to calibrate the chemical shift of the NMR spectra. Following chemical shift calibration, as well as baseline and phase corrections, the NMR spectra were displayed. Figure S14 presents the complete 1H NMR spectra measured after 2.5 hours, while Figure S16 provides the spectra obtained after 24 hours. Additionally, Figure S15 and Figure S17 offer zoomed-in views of the 5.4–6 ppm range, highlighting the H (C1’) doublets of pentose aminooxazoline used for yield comparisons.

### v3

@@ Conclusion
Our findings show that spin-controlled asymmetric processes due to the CISS effect can locally increase the coercivity of a ferromagnetic thin film. Contrary to classical expectations, temperature, an important environmental variable, actually enhances this process. Our experiments reveal that chirality-induced coercivity by RAO becomes stronger as the temperature increases. These results not only provide insights into the temperature dependence of CISS, an active area of research, but also reinforce the stability of spin-controlled mechanisms involving CISS across various environmental conditions. The inherent stability of CISS, which distinguishes it from classical magnetic phenomena, arises from spin-phonon couplings, which appear to be crucial for the high degree of spin polarization observed in CISS experiments. Thus, our results offer both a deeper understanding of CISS and highlight its significance in the context of the origin of life’s homochirality.
@@ Methods
Acknowledgments
We thank Naama Goren from the Hebrew University Applied Physics Department for her help in interpreting the Hall measurements and Maurice Saidian and Alexander Oginets from the Hebrew University Nano Center for Ni/Au evaporation. Y.P. acknowledges the funding from Marie Skłodowska-Curie Actions under Horizon Europe framework program (CISSE project No. 101071886), Carl Zeiss Stiftung (HYMMS project No. P2022-03-044), and U.S Air Force (grant No. FA8655-24-1-7390). Y.K. thanks the Israel Council of Higher Education’s VTT fellowship for women in STEM. S.F.O. and D.D.S. acknowledge the Harvard Origins of Life Initiative for funding and its members for fruitful discussions. S.F.O. further acknowledges support from the Kavli-Laukien fellowship program, Caltech startup funds, the William H. Hurt Scholarship Program, and the Fellows of King’s College Cambridge for their hospitality.
@@ 1.1 MOKE Setup
1.2 Effect of RAO Handedness
The MOKE experiment was repeated using both L- and D-enantiomers of RAO crystals. In Figure S2a, the magnetic hysteresis loops of Ni are presented for regions far from (top) and directly beneath (bottom) the D-RAO crystals. For the L-RAO sample, the magnetic properties of the bare Ni surface remained unchanged at all temperatures. In contrast, for the D-RAO sample, the coercive field of the bare Ni surface decreased at higher temperatures. Nonetheless, the difference in the coercive field between the regions beneath and far from the crystals increased with temperature, as shown in Figure S2b. In Figure S2c, this difference is compared for both L- and D-enantiomers, where a similar linear trend is observed for both. The chiral molecules act as strong pinning sites with a mechanism that does not favor a particular direction, as previously discussed in previous works [27, 17]. In addition, it was previously observed that a-chiral crystals such as glycine and salt do not cause a pinning effect on the surface (see Figure 3), and therefore, we attribute the pinning effects seen here to chirality-induced magnetization and conclude that it does not have a favored magnetization direction in this experimental setup.
@@ 1.7 MOKE Analysis Code
2 Hall Measurements
When an electric current flows through a conductor in a perpendicular magnetic field, the Lorentz force creates a transverse Hall voltage (). In ferromagnets, the anomalous Hall effect introduces an additional term from intrinsic magnetization. The total Hall voltage is: . Where is the regular Hall term and is the anomalous term due to the material’s magnetization. The anomalous Hall effect is commonly used to probe surface magnetization induced by chiral molecules [19, 8].
2.1 Hall Setup and Methods
Hall measurements were performed on three types of samples: Ni(40 nm)/Au(10 nm), Ni(7mn)/Au(7nm) patterned hall bar, and Ni(7nm)/Au(7nm)/Al2O3(1nm) patterned hall bar. Three types of molecules were studied: D-RAO, L-RAO, and a-chiral Glycine. The measurements were conducted under no magnetic field or a constant external magnetic field generated by an out-of-plane electromagnet powered by a Kepco BOP 100-4DL power supply. The current was supplied by Keysight B2901A precision source while perpendicular Hall voltage was measured in parallel by Keitley 2400 SourceMeter. Temperature was controlled by a Peltier element and was measured continuously using a Pt thermistor. The current was scanned from -1 mA to 1 mA and back and forth. These measurements were repeated 5 times for each direction and at different temperatures: 30°C, 40°C, 50°C, 60°C, 70°C, and 80°C.
@@ 2.2 Analysis of Hall Measurements
A linear relation was observed for each measurement of the Hall voltage as a function of the current. A linear function was fitted to each measurement, with the slope of the fit being the hall resistance,. For example, the five repetitions in the forward direction of the Hall IV curve and the corresponding fits for Ni/Au with RAO crystals at 50mT and 80°C are shown in Figure S8a. was averaged across all repetitions for each specific external field, temperature, and direction (forward or reverse), both with and without the D-RAO molecules. For each of these averages, a standard error was calculated. for all the temperatures and fields are represented in Figure S8. As shown, the behavior is similar for all magnetic fields, including those with negative values. We expected to observe a negative Hall voltage for the negative fields, but this was not the case. This, along with the relatively small external fields applied (up to 50 mT), suggests that the effect we measured is the Anomalous Hall Effect, and in the applied field range, the magnetization of the Ni did not change.
To observe the changes to after molecular deposition, we calculated the change in percentages as: The error is the standard error of mean between measurements.
2.3 Inconclusive Sign for Hall Effect Measurements
The magnetization induced by RAO crystals on a Si/Ni(40 nm)/Au(10 nm) thin film was measured using a simple Hall bar configuration, as shown in Figure S9a. Au pads were fabricated at the corners of the Si/Ni(40 nm)/Au(10 nm) film. A current () was applied, and the perpendicular Hall voltage () was measured under an external magnetic field of mT applied normal to the surface. Enantiopure D-RAO, L-RAO, and achiral glycine were drop-cast onto the substrate and left to dry. Hall voltages were measured before and after molecule adsorption to avoid variations due to device geometry or intrinsic magnetic properties between samples. Two samples were prepared with each RAO enantiomer and one with glycine. As shown in the optical images in Figure S9b, the crystals formed on the surface varied significantly between samples.
The Hall resistance () was extracted following the protocol described in the previous section and is presented in Figure S9c. In all chiral samples, changes significantly after molecule deposition, whereas the achiral sample shows no change. However, both the magnitude and sign of the change vary between samples of the same chirality. Moreover, the measurements do not exhibit a clear temperature dependence. Because of this variability, we cannot determine whether magnetization reversal occurs due to the molecules. This is likely because Au does not provide specific binding to RAO, leading to non-uniform molecular orientation both within and between samples. It has previously been shown that molecular orientation on a surface can invert preferential spin conductance [26]. The crystalline structure provides directional facets relative to the surface during nucleation, but these facets may differ between samples and across large crystal domains, as in this case.
2.4 Sequential Deposition
To address the variability observed in large-area samples, we performed a sequential deposition experiment using a fabricated Hall bar device. This geometry reduces sample-to-sample variation and ensures that crystal growth proceeds from existing nucleation sites, thereby preserving the same crystalline facets. The Hall bar was fabricated using standard photolithography with a Ni(7 nm)/Au(7 nm) channel and Au (50 nm) contact pads, as illustrated in Figure S10c. A current was applied through the channel and the transverse Hall voltage was measured, with the device shown in the optical image in Figure S10a.
Sequential deposition was carried out by drop-casting of 20 mM D-RAO solution onto the Hall bar. After each deposition, the Hall signal was recorded. In total, one, two, and four drops were deposited, with the final crystal morphology shown in Figure S10b. The resulting changes in Hall resistance () at C and C are shown in Figure S10d. The Hall signal increases systematically with each additional drop, although the overall magnitude of the effect remains small. As in the larger-area devices, no clear temperature dependence is observed; however, this method allows us to resolve a consistent directionality within a single sample under controlled chemical conditions.
2.5 Oxide Capping
To create a more directional surface chemistry, we evaporated a 1 nm Al2O3 capping layer onto a Hall bar fabricated as described in the previous section (Figure S11d). The Hall signal was measured before and after deposition of 40 mM D-RAO, L-RAO, and glycine solutions (optical images in Figure S11a–c). The RAO samples produced a more uniform crystal coverage under these conditions. To achieve this, the RAO solutions were supersaturated by dissolving 40 mM of D- or L-RAO in water at C until fully dissolved, followed by drop-casting onto the Hall bar while still hot. Crystals then formed upon cooling. Both D- and L-RAO showed more uniform crystallization compared to the uncapped case, whereas glycine did not.
The corresponding values are shown in Figure S11e. Both RAO enantiomers enhanced the Hall signal slightly and in opposite directions, while glycine also induced a noticeable change. As in previous measurements, no clear temperature dependence was observed. However, in this configuration, the chemical interactions are more controlled. The weak overall signal is attributed to the increased separation between the ferromagnetic layer and the molecules, since the additional oxide layer reduces interfacial coupling.
Overall, we are unable to conclude that the RAO molecules induced any net magnetization. The main limitation lies in the nonspecific surface chemistry of the RAO–Au interface, in contrast to more established systems such as thiol-functionalized polypeptides that form strong Au–S bonds, or RAO–magnetite interactions stabilized through hydrogen bonding. In our case, the RAO molecules are likely positioned farther from the substrate and lack directional binding, which reduces the possibility of generating a robust net magnetization. Thus, even if local magnetization changes occurred, they would not necessarily translate into a measurable exchange bias at the Ni/Au surface.
@@ 2.6 Hall Analysis Code
3 Linearity of the Energy Gap
The interaction given in Eq. (2) was derived in Ref.[11] and is based on electronically mediated interactions between localized spin moments and nuclear displacement. This formulation is in direct analogy with both indirect spin-spin interactions as well as the electronic contribution to the atomic force constant. The interaction tensor reflects the susceptibility associated with charge-spin fluctuations and is nonzero whenever the mediating electrons are either spin-polarized and/or spin-orbit coupled. In the present setup, both requirements can be presumed to be satisfied.
The close to linear temperature dependence of the dispersion relation is obtained from the Hartree approximation of the correction to the dispersion relation (see, e.g., Eq. (7) in Ref.[10]). The temperature-dependent factor in this correction can be written as
where is the thermal energy (with the Boltzmann constant and the temperature), is the magnetic moment per atom, is the unperturbed spin anisotropy, and is the high-energy cut-off for the magnon band.
For high temperatures (), the exponentials can, to a good approximation, be Taylor expanded to first order in . In this limit, the logarithm becomes independent of temperature, such that
Hence, in the parameter range relevant to the present context, the correction is close to linear for .